# Transformers Represent Belief State Geometry in their Residual Stream

### v1

@@ Abstract
What computational structure are we building into large language models when we train them on next-token prediction? Here, we present evidence that this structure is given by the meta-dynamics of belief updating over hidden states of the data-generating process. Leveraging the theory of optimal prediction, we anticipate and then find that belief states are linearly represented in the residual stream of transformers, even in cases where the predicted belief state geometry has highly nontrivial fractal structure. We investigate cases where the belief state geometry is represented in the final residual stream or distributed across the residual streams of multiple layers, providing a framework to explain these observations. Furthermore we demonstrate that the inferred belief states contain information about the entire future, beyond the local next-token prediction that the transformers are explicitly trained on. Our work provides a framework connecting the structure of training data to the computational structure and representations that transformers use to carry out their behavior.
@@ 1 Introduction
To test this framework, we conduct well-controlled experiments where we train transformers on data generated from processes with hidden ground truth structure, and then use our theory to make predictions about the geometry of internal activations. Even in cases where the framework predicts highly nontrivial fractal structure, our empirical results confirm these predictions (Figure 1). This provides a comprehensive explanation of how transformers encode and utilize information beyond the local next-token predictions they are explicitly trained on.
@@
In Section 2, we review the relevant theory from Computational Mechanics which motivates our prediction of the geometry of activations in the residual stream. Core to our work is the Mixed State Presentation; it describes the metadynamic of beliefs in the probability simplex over generator states. This formalism leads to a geometric structure that forms a natural hypothesis for the internal states of neural networks trained on prediction tasks.
@@ 2.1 Data generating processes
In this study, we assume that our training data is generated by an edge-emitting Hidden Markov Model (HMM)  [^1] For arbitrarily large, possibly non-ergodic HMMs with arbitrary initial distribution over the latent states, this does not limit the sophistication of training data.. Paths through the HMM produce sequences of tokens from a predefined vocabulary. Tokens ($x\in\mathcal{X}$) are emitted as we transition between hidden states ($s\in\mathcal{S}$). These transitions are governed by token-labeled transition matrices $\{T^{(x)}\}$, where each $T^{(x)}_{i,j}=\Pr(x,s_{j}|s_{i})$ represents the joint probability of emitting token $x$ and transitioning to state $s_{j}$, given that the model was in state $s_{i}$.
As a simple example, consider the HMM shown in Figure 2, that defines a process that generates strings of tokens of the form ...01R01R..., where R is a randomly generated 0 or 1. This HMM has 3 states: $\texttt{S}_{\texttt{0}}$, $\texttt{S}_{\texttt{1}}$, and $\texttt{S}_{\texttt{R}}$. Arrows of the form $s\xrightarrow{x:p\%}s^{\prime}$ denote the probability $\Pr(x,s^{\prime}|s)=p\%$ of moving to state $s^{\prime}$ and emitting the token $x$, given that the process was in state $s$.
Figure 2: An illustration of a Hidden Markov Model (HMM) and its components. The left side shows the HMM with states $\texttt{S}_{\texttt{0}}$, $\texttt{S}_{\texttt{1}}$, and $\texttt{S}_{\texttt{R}}$, and their respective transition probabilities. The right side displays the transition matrices $T^{(0)}$ and $T^{(1)}$ corresponding to token emissions 00 and $1$. Example training data is provided at the bottom, demonstrating a sequence generated by the HMM.
@@ 2.2 Mixed-state presentations
Computational Mechanics is a formalism that was developed in order to study the limits of prediction in chaotic systems, and has since expanded to a deep and rigorous theory of computational structure for any process [3]. One of its contributions is in providing a rigorous answer to what structures are necessary to perform optimal prediction [14]. Interestingly, Computational Mechanics shows that prediction is substantially more complicated than generation. Informally, the reason for this is that even if one knows the underlying hidden generative structure, one must still perform computational work in order to infer which state the generative process is in given finite observations of data. Thus, there are conceptually two distinct inference processes related to prediction: learning the hidden structure of the data generating process, and subsequently inferring which state the data generating process is in given finite observations of data.
Computational Mechanics formally captures the computational structure of the latter inference process with the Mixed-State Presentation (MSP). Whereas the HMM of a data generating process describes a generative structure (Figure 3A), the Mixed-State Presentation (MSP) is the structure of prediction (Figure 3B). The MSP answers the question of how an optimal observer updates their beliefs over the states of the data-generating process given finite observations of tokens [12, 8]. If the observer is in a belief state given by a probability distribution $\boldsymbol{\eta}$ (a row vector) over the hidden states of the data generating process, then the update rule for the new belief state $\boldsymbol{\eta}\textquoteright$ ’ given that the observer sees a new token $x$ is:
@@
This process of belief updating is itself another HMM, where hidden states are associated with belief states, and paths leading to particular belief states are the observed token sequence. The probability of transitioning from belief states $\boldsymbol{\eta}$ to $\boldsymbol{\eta}^{\prime}$ of Eq. (1) is given by $\Pr(\boldsymbol{\eta}^{\prime},x|\boldsymbol{\eta})=\boldsymbol{\eta}T^{(x)}\boldsymbol{1}$.
For stationary processes, the optimal initial belief state is given by the stationary distribution $\boldsymbol{\eta}_{\varnothing}=\boldsymbol{\pi}$ over latent states (the left-eigenvector of the transition matrix $T=\sum_{x}T^{(x)}$ associated with the eigenvalue of 1).
Figure 3: The generative structure (Left) implies a unique minimal metadynamic over belief states as a predictor synchronizes to the hidden state of the world as it observes more context (Middle). These belief states are embedded in a probability simplex over the minimal generative states (Right). Transformers distinguish these belief states, even if they imply the same probability distribution over the next-token (as with $\boldsymbol{\eta}_{10}$ and $\boldsymbol{\eta}_{01}$).
Each belief state of the MSP is a probability distribution over the states of the data generating process. We can plot these in a probability simplex, as indicated in Figure 3C. Note that each belief state—as a distribution over generator states—generates a probability density over all possible futures. Distributions over the entire future clearly carry more than just next-token information. We can see this in our example in Figure 2. The states $\boldsymbol{\pi}$, $\boldsymbol{\eta}_{10}$, and $\boldsymbol{\eta}_{01}$ all have the same next-token prediction despite being distinct belief states.
@@ 2.3 Finding the belief simplex in the residual stream
We begin by training a transformer on data generated by a ground-truth HMM. We consider activations of the residual stream (in either a single layer, or a concatenation of layers) induced by all possible input sequences (Figure 4AB), and at all context window positions. Thus our dataset consists of a set of activations, $a\in\mathbb{R}^{d_{\text{resid}}}$ when we consider a single layer at a given position, where $d_{\text{resid}}$ is the residual stream dimension. For each input sequence, our framework tells us which belief state the input corresponds to, and the probability distribution over hidden states of the generative process that corresponds to this belief state. These probability distributions can be thought of as points $b\in\mathbb{R}^{|\mathcal{S}|}$, where $|\mathcal{S}|$ denotes the number of hidden states in the generative process.
@@ 3.1 Belief state geometry is linearly represented in the residual stream
To test these predictions, we analyzed the final layer of the transformer’s residual stream, before the layer norm and unembedding. Using linear regression, we identified a 2D subspace of the 64-dimensional residual activations that best matched the ground-truth belief distributions from the MSP. Remarkably, the geometry of this 2D subspace closely resembled the predicted fractal structure (Figure 5C), providing strong evidence that the transformer had indeed learned to represent the geometry of the belief states in its residual stream.
We ran a number of controls, shown in Figure 6 to make sure these results were not artifacts. First, we performed our analysis at multiple points through training and found that the simplex structure emerged gradually (Figure 6A), suggesting that the detailed fractal structure found at the end of training was not a trivial consequence of the model architecture or initialization. Quantifying the mean squared error of the regression revealed a decrease in the errors of belief-state geometry representation throughout the course of training (Figure 6D, first 4 bars). We also performed cross-validation, where only 20% of all input-activation pairs were used to train the regression, and then the held out 80% data was used to visualize and analyze the result, and found that the geometry of the belief state was still well represented (Figure 6B) and that the mean squared error was similar to that of the full regression (Figure 6D, red vs. purple bars). Finally, we ran a control that preserved the fractal geometry in the simplex but shuffled the input-point correspondences, resulting in the regression mapping all points to the simplex center due to the lack of discoverable structure (Figure 6C).
These results provide compelling evidence for our central claim: transformers trained on data with hidden generative structure will learn to represent the geometry of belief states in their residual stream. The close match between the predicted fractal structure and the empirical geometry of the residual activations, even for the highly complex Mess3 Process, suggests that this geometry is a fundamental aspect of how transformers build predictive representations of their input data.
@@ 3.2 Belief state geometry represents information beyond next-token prediction
The Random–Random–XOR (RRXOR) process, shown in Figure 7A, has these types of degeneracies [13]. The RRXOR process has 36 distinct belief states, and can be geometrically represented in a 4-simplex. In Figure 7B we visualize the simplex geometry by projecting down to 2 dimensions. After training, we run our regression analysis to see if the belief state geometry is represented in the residual stream. Unlike our previous results for the Mess3 process, we find that the belief state geometry is not represented in the final residual stream before the unembedding. However, when we take the residual stream activations of all layers and concatenate them, we find a veridical representation (Figure 7C).
@@
Next, for every pairwise distances of belief state representations, we compute the corresponding distance in the ground truth next token probabilities. The scatter plots on the right of Figure 7D compares these distances. For the RRXOR process, the correlation is low ($R^{2}=0.31$), indicating that the belief state representations in the residual stream is not captured by the next-token predictions. For the Mess3 process, the pairwise structure of next-token predictions is preserved in the belief state representation. This difference between the RRXOR and Mess3 results is expected from the theory. Since the RRXOR process contains degeneracies in its belief states with respect to next-token predictions, the discovered belief state geometry is not well-explained by the next token predictions.
@@ 4.1 Belief state geometry: Why?
During training, models do not directly see the hidden structure generating the data—they only see data. So how is it that the pretrained model infers belief states in the simplex of the generator states? In fact, there are infinitely many distinct HMMs that can generate the same stochastic process [4]. Despite this fact, any stochastic process has a canonical vector-space representation in the space of probability densities over all possible future token sequences [16]. This corresponds to a minimal HMM representation, and we should expect that transformers learn belief state geometry in the simplex over these minimal generator states. This canonical geometry for a process provides an explanation for why we were able to find the belief state geometry represented in our transformers.
In the jargon of Computational Mechanics, it is notable that the recurrent belief states map onto the causal states of the process’ $\epsilon$-machine [3]. The recurrent states must be distinguished for prediction, even if there exists a generative HMM that is smaller than the $\epsilon$-machine. Typically, data can be generated by a much smaller mechanism than is required to predict it. Accordingly, transformers and other pretrained models learn much more than just an accurate generative model of the world—they also learn how to synchronize to the hidden state of the world through observed context.
We expect our main results do not depend strongly on the particular neural network architecture of transformers, except that they (i) utilize a residual stream and (ii) were pretrained on future-token prediction. Indeed, we expect to find the same belief state geometry in other neural network architectures like Mamba [6], although this has not yet been verified. The generality of our results is thus a powerful architecture-independent insight for interpretability. Pretraining will generally induce mixed-state geometry. We should then be able to work backwards from this knowledge and the knowledge of the neural network architecture to determine both its world model and how it performs Bayesian updating over its hidden states.
Corollary 2 of [14] implies that, for a recurrent neural network to perform as well as possible on next token prediction, all information about the past necessary to predict the entire future as well as possible must be present in the latent state of the network. The implication for feedforward networks like transformer architectures is much less obvious, but it seems that the analogous condition is that: To perform as well as possible on next-token prediction requires that all information about the past necessary to predict the entire future as well as possible must be present across the layers of each position’s residual stream. Our results here support this interpretation.
@@ 4.3 What are features?
4.4 Limitations
Our experiments were conducted on small HMMs with 3 states and vocabularies of size 2 or 3—although both RRX and Mess3 express complexity via infinite Markov order. In practice, real-world LLMs are much larger, with vocabularies exceeding 50,000 tokens and more sophisticated data-generating structures. As such, the applicability to larger, more complex settings remains uncertain.
This work focused on ergodic and stationary processes. Computational mechanics can address both non-stationary and non-ergodic processes, which are particularly relevant to in-context learning, but we have not yet explored these extensions empirically in LLMs.
We do not yet know if these LLMs develop something like a circuit for implementing Bayesian updates, as would be natural for MSP dynamics. So far, we have only established the representation of belief states and their geometry. It is also unclear how these LLMs generalize when driven by a different process.
Computational Mechanics is primarily concerned with optimal prediction, but LLMs in practice are not perfectly optimal. Further work is needed to understand how near-optimality and non-optimality influence belief state representations. Progress in this direction could offer insights into evolving structures during training.

### v2

@@ Abstract
What computational structure are we building into large language models when we train them on next-token prediction? Here, we present evidence that this structure is given by the meta-dynamics of belief updating over hidden states of the data-generating process. Leveraging the theory of optimal prediction, we anticipate and then find that belief states are linearly represented in the residual stream of transformers, even in cases where the predicted belief state geometry has highly nontrivial fractal structure. We investigate cases where the belief state geometry is represented in the final residual stream or distributed across the residual streams of multiple layers, providing a framework to explain these observations. Furthermore we demonstrate that the inferred belief states contain information about the entire future, beyond the local next-token prediction that the transformers are explicitly trained on. Our work provides a general framework connecting the structure of training data to the geometric structure of activations inside transformers.
@@ 1 Introduction
To test this framework, we conduct well-controlled experiments where we train transformers on data generated from processes with hidden ground truth structure, and then use our theory to make predictions about the geometry of internal activations. Even in cases where the framework predicts highly nontrivial fractal structure, our empirical results confirm these predictions (Figure 1). This provides a comprehensive explanation of how transformers encode information beyond the local next-token predictions they are explicitly trained on.
@@
In Section 2, we review the relevant theory from computational mechanics which motivates our prediction of the geometry of activations in the residual stream. Core to our work is the mixed-state presentation, which describes the metadynamic of beliefs in the probability simplex over states of the data generating process. This formalism leads to a geometric structure that forms a natural hypothesis for the internal states of neural networks trained on prediction tasks.
@@ 2.1 Data generating processes
In this study, we assume that our training data is generated by an edge-emitting Hidden Markov Model (HMM) [^1] For arbitrarily large, possibly non-ergodic HMMs with arbitrary initial distribution over the latent states, this does not limit the sophistication of training data.. Paths through the HMM produce sequences of tokens from a predefined vocabulary. Tokens ($x\in\mathcal{X}$) are emitted as we transition between hidden states ($s\in\mathcal{S}$). These transitions are governed by token-labeled transition matrices $\{T^{(x)}\}$, where each $T^{(x)}_{i,j}=\Pr(x,s_{j}|s_{i})$ represents the joint probability of emitting token $x$ and transitioning to state $s_{j}$, given that the HMM was in state $s_{i}$.
As a simple example, consider the HMM shown in Figure 2, that we call the Zero-One-Random (Z1R) process. The Z1R process generates strings of tokens of the form ...01R01R..., where R is a randomly generated 0 or 1. This HMM has 3 states: $\texttt{S}_{\texttt{0}}$, $\texttt{S}_{\texttt{1}}$, and $\texttt{S}_{\texttt{R}}$. Arrows of the form $s\xrightarrow{x:p\%}s^{\prime}$ denote the probability $\Pr(x,s^{\prime}|s)=p\%$ of moving to state $s^{\prime}$ and emitting the token $x$, given that the process was in state $s$.
Figure 2: An illustration of a Hidden Markov Model (HMM) and its components. The left side shows the HMM with states $\texttt{S}_{\texttt{0}}$, $\texttt{S}_{\texttt{1}}$, and $\texttt{S}_{\texttt{R}}$, and their respective transition probabilities. The right side displays the transition matrices $T^{(\texttt{0})}$ and $T^{(\texttt{1})}$ corresponding to token emissions 0 and 1. Example training data is provided at the bottom, demonstrating a sequence generated by the HMM.
@@ 2.2 Mixed-state presentations
Computational mechanics is a formalism that was developed in order to study the limits of prediction in chaotic systems, and has since expanded to a deep and rigorous theory of computational structure for any process [4]. One of its contributions is in providing a rigorous answer to what structures are necessary to perform optimal prediction [31].
Interestingly, computational mechanics shows that prediction is substantially more complicated than generation [19, 30]. Informally, the reason for this is that even if an agent knows the underlying hidden generative structure that creates the data they are trying to predict, the agent must still perform computational work in order to infer which state the generative process is in, given finite observations of data. Thus, there are conceptually two distinct inference processes related to prediction: learning the hidden structure of the data generating process, and subsequently inferring which state the data generating process is in given finite observations of data.
Computational mechanics formally captures the computational structure of the latter inference process with the mixed-state presentation (MSP). Whereas the HMM of a data generating process describes a generative structure (Figure 3A), the MSP is the structure of prediction (Figure 3B). The MSP answers the question of how an optimal observer updates their beliefs over the states of the data-generating process given finite observations of tokens [27, 15]. If the observer is in a belief state given by a probability distribution $\boldsymbol{\eta}$ (a row vector) over the hidden states of the data generating process, then the update rule for the new belief state $\boldsymbol{\eta}\textquoteright$ ’ given that the observer sees a new token $x$ is:
@@
This process of belief updating is itself another HMM, where hidden states are associated with belief states, and paths leading to particular belief states are the observed token sequence. The probability of transitioning from belief state $\boldsymbol{\eta}$ to $\boldsymbol{\eta}^{\prime}=\boldsymbol{\eta}T^{(x)}/\boldsymbol{\eta}T^{(x)}\boldsymbol{1}$ of Eq. (1) is simply given by the probability $\boldsymbol{\eta}T^{(x)}\boldsymbol{1}$ of observing an $x$ from $\boldsymbol{\eta}$. For stationary processes, the optimal initial belief state is given by the stationary distribution $\boldsymbol{\eta}_{\varnothing}=\boldsymbol{\pi}$ over latent states (the left-eigenvector of the transition matrix $T=\sum_{x}T^{(x)}$ associated with the eigenvalue of 1).
Figure 3: (A) An example generative structure called the zero-one-random process (Z1R), since it generates data of the form ...01R01R01R... where R is a random bit. (B) The generative structure implies a unique metadynamic over belief states as a predictor synchronizes to the hidden state of the world as it observes more context. This predictive structure is called the mixed-state presentation (MSP). We label belief states with $\boldsymbol{\eta}_{w}$ where $w$ is the shortest string of emissions which leads to that belief state. (C) Belief states are distributions over generator states and can be embedded in a probability simplex. To read off this distribution for a given belief state, $\boldsymbol{\eta}$, one measures the perpendicular distance from the point to each edge of the simplex, shown as blue, green, and red lines in this example. These distances directly give the probabilities for each state. Thus, the vertices represent states of certainty over one generator state, since the perpendicular distance is nonzero to only one of the edges of the simplex. (D) Plotting the belief state distributions in the probability simplex gives the belief state geometry.
Each belief state of the MSP is a probability distribution over the states of the data generating process. Belief states thus inherit a natural geometry; we can plot these belief states in a probability simplex, as indicated in Figure 3C. We refer to the geometric arrangement of the points on the simplex as the belief state geometry. Note that each belief state—as a distribution over generator states—induces a probability density over all possible futures. Distributions over the entire future clearly carry more than just next-token information. We can see this in our example in Figure 3D. The states $\boldsymbol{\pi}$, $\boldsymbol{\eta}_{\texttt{10}}$, and $\boldsymbol{\eta}_{\texttt{01}}$ all have the same next-token prediction despite being distinct belief states.
@@ 2.3 Finding the belief simplex in the residual stream
We begin by training a transformer on data generated by a ground-truth HMM. We consider activations of the residual stream (in either a single layer, or a concatenation of layers) induced by all possible input sequences (Figure 4AB), and at all context window positions. Thus our dataset consists of a set of activations, $a\in\mathbb{R}^{d_{\text{resid}}}$, when we consider a single layer at a given position, where $d_{\text{resid}}$ is the residual stream dimension. For each input sequence, our framework tells us which belief state the input corresponds to, and the probability distribution over hidden states of the generative process that corresponds to this belief state. These probability distributions can be thought of as points $b\in\mathbb{R}^{|\mathcal{S}|}$, where $|\mathcal{S}|$ denotes the number of hidden states in the generative process.
@@ 3.1 Belief state geometry is linearly represented in the residual stream
To test these predictions, we analyzed the final layer of the transformer’s residual stream, before the layer norm and unembedding. Using linear regression, we identified a 2D subspace of the 64-dimensional residual activations that best matched the ground-truth belief distributions from the MSP. Remarkably, the geometry of this 2D subspace closely resembled the predicted fractal structure (Figure 5C), providing strong evidence that the transformer had indeed learned to represent the geometry of the belief states in its residual stream [^3] These results hold when analyzing the activations after the final layer norm as well, see Figure S1..
We ran a number of controls, shown in Figure 6, to make sure these results were not artifacts. First, we performed our analysis at multiple points through training and found that the simplex structure emerged gradually (Figure 6A), suggesting that the detailed fractal structure found at the end of training was not a trivial consequence of the model architecture or initialization. Second, quantifying the mean squared error of the regression revealed a decrease in the errors of belief-state geometry representation throughout the course of training (Figure 6D, first 4 bars). In addition, more faithful representation of the belief state geometry in the residual stream corresponded to lower cross entropy loss (Figure S2). Third, we performed cross-validation, where only 20% of all input-activation pairs were used to train the regression, and then the held out 80% data was used to visualize and analyze the result, and found that the geometry of the belief state was still well represented (Figure 6B) and that the mean squared error was similar to that of the full regression (Figure 6D, red vs. purple bars). Finally, we ran a control that preserved the fractal geometry in the simplex but shuffled the input-point correspondences, resulting in the regression mapping all points to the simplex center due to the lack of discoverable structure (Figure 6C).
These results provide compelling evidence for our central claim: transformers trained on data with hidden generative structure will learn to represent the geometry of belief states in their residual stream. The close match between the predicted fractal structure and the empirical geometry of the residual activations, even for the highly complex Mess3 MSP, suggests that this geometry is a fundamental aspect of how transformers build predictive representations of their input data.
@@ 3.2 Belief state geometry represents information beyond next-token prediction
The Random–Random–XOR (RRXOR) process, shown in Figure 7A, has these types of degeneracies [28]. The MSP of the RRXOR process has 36 distinct belief states, and can be geometrically represented in a 4-simplex. In Figure 7B we visualize the simplex geometry by projecting down to 2 dimensions. After training, we run our regression analysis to see if the belief state geometry is represented in the residual stream. Unlike our previous results for the Mess3 process, we find that the belief state geometry is not represented in the final residual stream before the unembedding. However, when we take the residual stream activations of all layers and concatenate them, we find a veridical representation (Figure 7C).
@@
Next, for every pairwise distances of belief state representations, we compute the corresponding distance in the ground truth next token probabilities. The scatter plots on the right of Figure 7D compares these distances. For the RRXOR process, the correlation with next-token predictions is relatively low ($R^{2}=0.31$) compared to the correlation with the belief state geometry ($R^{2}=0.95$), indicating that belief state geometry captures structure in the residual stream not captured by next-token predictions. For the Mess3 process, the pairwise structure of next-token predictions is preserved in the belief state representation. This difference between the RRXOR and Mess3 results is expected from the theory. Since the RRXOR process contains distinct belief states with the same optimal next-token distribution, the discovered belief state geometry is not well-explained by the next token predictions.
@@ 4.1 Belief state geometry: Why?
During training, models do not directly see the hidden structure generating the data—they only see data. So how is it that the pretrained model infers belief states in the simplex of the generator states? In fact, there are infinitely many distinct HMMs that can generate the same stochastic process [5]. Despite this, any stochastic process has a canonical vector-space representation in the space of probability densities over all possible future token sequences [33]. This corresponds to a minimal HMM representation, and we should expect that transformers learn belief state geometry in the simplex over these minimal generator states. This canonical geometry for a process provides an explanation for why we were able to find the belief state geometry represented in our transformers.
In the jargon of computational mechanics, it is notable that the recurrent belief states map onto the causal states of the process’ $\epsilon$-machine [4]. The recurrent states must be distinguished for prediction, even if there exists a generative HMM that is smaller than the $\epsilon$-machine. Typically, data can be generated by a much smaller mechanism than is required to predict it [19, 30]. Accordingly, transformers and other pretrained models learn much more than just an accurate generative model of the world—they also learn how to synchronize to the hidden state of the world through observed context.
We expect our main results do not depend strongly on the particular neural network architecture of transformers, except that they (i) utilize a residual stream and (ii) were pretrained on future-token prediction. Indeed, we expect to find the same belief state geometry in other neural network architectures like Mamba [12], and recent work building on our discovery has shown that recurrent neural networks do represent the belief state geometry for the processes tested in our experiments [26]. The generality of our results is thus a powerful architecture-independent insight for interpretability. Pretraining will generally induce mixed-state geometry. We should then be able to work backwards from this knowledge and the knowledge of the neural network architecture to determine both its world model and how it performs Bayesian updating over its hidden states.
Corollary 2 of [31] implies that, for a recurrent neural network to perform as well as possible on next token prediction, all information about the past necessary to predict the entire future as well as possible must be present in the latent state of the network. The implication for feedforward networks like transformer architectures is much less obvious, but it motivates the hypothesis that to perform as well as possible on next-token prediction requires that all information about the past necessary to predict the entire future as well as possible must be present across the layers of each context window position’s residual stream. Our results here support this interpretation.
@@
It is also interesting to consider the linearity of the belief state representations we found, and if the linearlity is strictly necessary for a network to optimize next-token cross-entropy loss. In fact, recent work has constructed a transformer by hand (i.e. a human being selected the weights) that perfectly solves the RRXOR task, but where the belief state geometry is not linearly represented [11]. It is thus a non-trivial empirical finding that transformer architectures trained via stochastic gradient descent find solutions in which the belief state geometry is linearly represented.
@@ 4.3 What are features?
4.4 Limitations and applicability to more realistic settings
Our experimental validation focused on small-scale systems, using HMMs with only 3-5 states and vocabularies of 2-3 tokens. While these systems exhibited meaningful complexity through infinite Markov order in both the RRXOR and Mess3 processes, they represent a significant simplification compared to the sophisticated architectures and massive vocabularies (>50,000 tokens) of modern language models trained on natural language data. This scale limitation raises important questions about how our framework extends to larger and more complex settings.
For realistic systems like board games or natural language, the dimensionality of the belief state simplex would exceed the dimension of the residual stream. This suggests that if transformers represent belief state geometry in these domains, they must do so in a compressed form. Understanding the nature of this compression—what information is preserved and what is discarded—represents a crucial direction for future work.
As the size of the minimal generating structure and the context window length grows, explicitly computing ground-truth belief states becomes intractable. We will need new methodologies for validating our framework in settings where we cannot directly compute the MSP structure.
We limited ourselves to the simpler stationary ergodic setting in this paper to empirically verify important basic features of transformer representations. However, the framework presented in this work will naturally extend to non-stationary [29] and non-ergodic processes [6], which is important for the extension to real-world tasks. For example, in in-context learning the underlying process is inherently nonergodic. In the more general setting, we also expect transformers to represent belief-state geometry, although that geometry will then reflect the more general types of non-stationary and non-ergodic processes representative of real-world data like natural language [7, 8].
In the experiments we ran, we generated training data using HMMs. A natural question is what to expect with training data generated from non-HMM sources. It is important to note that any dataset made of sequences of tokens used to train a transformer can be represented as being generated from an HMM, if one allows for non-ergodicity, non-stationarity, and infinite states. For example, in the case of training data that takes the form of parenthesis matched strings, one would usually conceive of this as being generated by a push-down automaton. However, this can equivalentaly be represented by an HMM consisting of an infinite chain of hidden states. Another interesting point that we did not focus on in our presentation for the sake of simplicity, is that even if one does not assume an HMM for the generating machine, the question of the computational structure of an optimal predictor is naturally an HMM, with hidden states given by distinct distributions over the future (whether the generator is an HMM or a Turing Machine, or something else) [33].
In this work we verified that belief states are linearly represented in the residual stream, even in cases where the belief states form intricate fractal geometries. Although this is quite suggestive that these representations are used to implement some form of Bayesian updating, we did not directly test this. In other words, we do not yet know if these LLMs develop something like a circuit for implementing Bayesian updates, as would be natural for MSP dynamics. So far, we have only established the representation of belief states and their geometry.
Computational mechanics is primarily (but not completely, see Ref. [21]) concerned with optimal prediction, but LLMs in practice are not perfectly optimal. Further work is needed to understand how near-optimality and non-optimality influence belief state representations. Progress in this direction could offer insights into evolving structures during training.
Moving beyond stationary ergodic HMM training data, we still generically expect fractal-like structure in the residual-stream activations. Why? The fractal structure is a result of Bayesian updating applied to the “non-deterministic” computational structure of HMMs [16]. With stationary processes, this same Bayesian map is folded into itself time and again, producing a very clean fractal. For natural language, even in the absence of stationarity, there will nevertheless be resonances of non-deterministic computational structure, which would imply a type of folding beliefs in a self-similar manner. However, a more rigorous version of these statements and the empirical validation is left for future work.